# Simultaneous Source Separation, Synchronization, Localization and Mapping for 6G Systems

### v1

@@ IV-D Data Association
Data assocation between measurements and PVAs is described in a redundant formulation by the PVA-oriented (and PA-oriented) association vector $\underline{\bm{\mathsfbr{a}}}_{n}^{(i)}\triangleq[\underline{\mathsfbr{a}}_{1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\underline{\mathsfbr{a}}_{J,n}^{(i)}\hskip 2.84526pt\underline{\mathsfbr{a}}_{J+1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\underline{\mathsfbr{a}}_{J+\mathsfbr{K}^{(i)}_{n-1},n}^{(i)}]^{\mathrm{T}}$ and by the measurement-oriented association vector $\overline{\bm{\mathsfbr{a}}}_{n}^{(i)}\triangleq[\overline{\mathsfbr{a}}_{1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\overline{\mathsfbr{a}}_{{\mathsfbr{M}}_{n}^{(i)},n}^{(i)}]^{\mathrm{T}}$. See [25, 23, 6, 17] for details.
@@ V-1 Joint Posterior PDF and Factor Graph
Using Bayes’ rule and independence assumptions related to the state-transition PDFs, the prior PDFs, and the likelihood model (for details please see [23, 6, 19]), and for fixed (observed) measurements $\bm{z}_{1:n}$ (the numbers of measurements ${M}^{(i)}_{n}$ are fixed and not random anymore) the joint posterior PDF of $\bm{\mathsfbr{y}}_{0:n}$, $\bm{\mathsfbr{x}}_{0:n}$, $\underline{\bm{\mathsfbr{a}}}_{1:n}$, and $\overline{\bm{\mathsfbr{a}}}_{1:n}$, conditioned on $\bm{z}_{1:n}$ is given by (15), where the individual LHF-related factors are respectively $\underline{q}_{\text{BS}}\big(\underline{a}^{(i)}_{j,n},\bm{x}_{n}^{(i)},{b}^{(j)}_{\text{bs},n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$, $\underline{q}\big(\underline{\bm{y}}^{(i)}_{k,n},\bm{x}_{n}^{(i)},{b}^{(\underline{j}(k))}_{\text{bs},n}\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt,\underline{a}^{(i)}_{k+J,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$ with $\underline{j}(k)\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt(k\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt-\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt1)\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\mod\hskip-0.85358pt\hskip-0.85358ptJ+1$ (see (9)), and $\overline{q}\big(\overline{\bm{y}}^{(i)}_{m,n},\bm{x}^{(i)}_{n},\bm{b}_{\text{bs},n},\overline{j}^{(i)}_{m,n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$ that will be discussed next.
@@
$\underline{q}_{\text{BS}}\big(\bm{x}_{n}^{(i)},{b}^{(j)}_{\text{bs},n},\underline{a}^{(i)}_{j,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$
$\hskip 28.45274pt\triangleq\begin{cases}\displaystyle\hskip 0.85358pt\frac{p_{\mathrm{d}}f\big({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\big|\hskip 0.85358pt\bm{x}^{(i)}_{n},{b}^{(j)}_{\text{bs},n}\big)}{\mu_{\mathrm{fp}}\hskip 0.85358ptf_{\mathrm{fp}}\big({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\big)}\hskip 0.85358pt,&\!\!\hskip-0.85358pt\underline{a}^{(i)}_{j,n}\!=\hskip-0.85358ptm\\[9.95845pt] 1\!-\hskip-0.85358ptp_{\mathrm{d}}\hskip 0.85358pt,&\!\!\hskip-0.85358pt\underline{a}^{(i)}_{j,n}\!=\hskip-0.85358pt0\end{cases}$ (16)
@@
$\underline{q}\big(\underline{\bm{p}}^{(i)}_{k,\text{va}},\underline{r}^{(i)}_{k,n}=1,\bm{x}_{n}^{(i)},{b}^{(\underline{j}(k))}_{\text{bs},n},\underline{a}^{(i)}_{k,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$
$\triangleq\begin{cases}\displaystyle\hskip 0.85358pt\frac{p_{\mathrm{d}}f\big({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\big|\hskip 0.85358pt\bm{x}^{(i)}_{n},\underline{\bm{p}}^{(i)}_{k,\text{va}}\hskip-0.85358pt,{b}^{(\underline{j}(k))}_{\text{bs},n}\big)}{\mu_{\mathrm{fp}}\hskip 0.85358ptf_{\mathrm{fp}}\big({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\big)}\hskip 0.85358pt,&\!\!\hskip-0.85358pt\underline{a}^{(i)}_{k+J,n}\!=\hskip-0.85358ptm\\[9.95845pt] 1\!-\hskip-0.85358ptp_{\mathrm{d}}\hskip 0.85358pt,&\!\!\hskip-0.85358pt\underline{a}^{(i)}_{k+J,n}\!=\hskip-0.85358pt0\end{cases}$ (17)
and $\underline{q}\big(\underline{\bm{x}}^{(i)}_{k,n},\underline{r}^{(i)}_{k,n}=0,\bm{x}_{n}^{(i)},\underline{a}^{(i)}_{k,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)\hskip-0.85358pt\triangleq\hskip-0.85358pt{\delta_{\underline{a}^{(i)}_{k,n}}}$.
The pseudo LHF related to new PVA $m$ and MT $i$ is given by
$\overline{q}\big(\overline{\bm{y}}^{(i)}_{m,n},\bm{x}_{n}^{(i)},\bm{b}_{\text{bs},n},\overline{j}^{(i)}_{m,n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$
$\hskip 14.22636pt\triangleq\begin{cases}0\hskip 0.85358pt,&\hskip-2.84526ptj\!\neq\hskip-0.85358pt\overline{j}^{(i)}_{m,n}\\[2.84526pt] \overline{q}\big(\overline{\bm{y}}^{(i)}_{m,n},\bm{x}_{n}^{(i)},{b}^{(j)}_{\text{bs},n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)\hskip 0.85358pt,&\hskip-2.84526ptj\!=\hskip-0.85358pt\overline{j}^{(i)}_{m,n}\end{cases}$ (18)
@@
$\hskip-2.84526pt\overline{q}\big(\overline{\bm{p}}^{(j,i)}_{\text{va},m},\overline{r}^{(j,i)}_{m,n}=1,\bm{x}_{n}^{(i)},{b}^{(j)}_{\text{bs},n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$
$\hskip-2.84526pt\triangleq\begin{cases}0\hskip 0.85358pt,&\hskip-2.84526pt\overline{a}^{(i)}_{m,n}\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358ptk\\[2.84526pt] \frac{\mu_{\mathrm{n}}f_{\mathrm{n}}(\overline{\bm{p}}^{(j,i)}_{\text{va},m}\hskip 0.85358pt|\hskip 0.85358pt\bm{x}_{n}^{(i)})f({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\hskip 0.85358pt|\bm{x}_{n}^{(i)},\overline{\bm{p}}^{(j,i)}_{\text{va},m},{b}^{(j)}_{\text{bs},n})}{J\,\mu_{\mathrm{fp}}f_{\mathrm{fp}}({\bm{z}_{\text{bs}}}^{(i)}_{m,n})}\hskip 0.85358pt,&\hskip-2.84526pt\overline{a}^{(i)}_{m,n}\hskip-0.85358pt=\hskip-0.85358pt0\end{cases}$ (19)
and $\overline{q}\big(\overline{\bm{p}}^{(j,i)}_{\text{va},m},\overline{r}^{(j,i)}_{m,n}=0,\overline{a}^{(i)}_{m,n},\bm{x}_{n}^{(i)},{b}^{(j)}_{\text{bs},n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)\hskip-0.85358pt\triangleq\hskip-0.85358ptf_{\text{d}}\big(\overline{\bm{p}}^{(j,i)}_{\text{va},m}\big)$, respectively. Finally, the binary indicator functions that check consistency for any pair $(\underline{a}^{(i)}_{k,n},\overline{a}^{(i)}_{m,n})$ of PVA/PA-oriented and measurement-oriented association variable at time $n$,
@@
$\hskip-5.69054pt\triangleq\begin{cases}0,&\underline{a}^{(i)}_{k,n}=m,\overline{a}^{(i)}_{m,n})\neq k\text{ or }\underline{a}^{(i)}_{k,n}\neq m,\overline{a}^{(i)}_{m,n})=k\\[0.28453pt] 1\hskip 0.85358pt,&\text{otherwise}.\end{cases}$ (20)
@@ V-2 Confirmation of PVAs and State Estimation
We estimate the MT states $\bm{x}^{(i)}_{n}$ (including ${b}_{\text{mt}\,n}^{(i)}$), the BS biases ${b}_{\text{bs}\,n}^{(j)}$, and the PVA states $\bm{\mathsfbr{p}}_{\text{va},k}$ from all measurements $\bm{z}_{1:n}$ up to time $n$ using the respective minimum mean-square error (MMSE) estimates [26, Ch. 4] $\hat{{\bm{x}}}^{(i)}_{n}=[\hat{\tilde{\bm{x}}}^{(i)}_{n}\hskip 2.84526pt\hat{\bm{v}}^{(i)}_{n}\hskip 2.84526pt\hat{b}_{\text{mt},n}^{(i)}]$ and $\hat{b}_{\text{bs}\,n}^{(j)}$, which are calculated by the expected values of the respective marginal posterior distributions $f(\bm{x}^{(i)}_{n}|\bm{z}_{1:n})$, $f({b}_{\text{bs},n}^{(j)}|\bm{z}_{1:n})$. Detection of the PVAs $k\!\in\!\{1,\dots,K_{n}\}$ relies on the posterior existence probabilities $p(r_{k,n}\!=\!1|\bm{z}_{1:n})=\int f(\bm{p}_{\text{va},k},r_{k,n}\!=\!1|\bm{z}_{1:n})\,\mathrm{d}\bm{p}_{\text{va},k}$ and the conditional posteriors $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})=f(\bm{p}_{\text{va},k},r_{k,n}\!=\!1|\bm{z}_{1:n})/p(r_{k,n}\!=\!1|\bm{z}_{1:n})$. A PVA is confirmed if $p(r_{k,n}\!=\!1|\bm{z}_{1:n})>p_{\mathrm{cf}}$ [27], and pruned if $p(r_{k,n}\!=\!1|\bm{z}_{1:n})<p_{\text{pr}}$, except for k=1k\!=\!1 of each BS, which is always retained. For existing PVAs, we determine the MMSE estimate $\hat{\bm{p}}_{\text{va},k}$ by calculating the expected value of $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})$. Since direct marginalization of (15) is intractable, sequential message passing via the SPA on a factor graph [28] provides efficient approximations (“beliefs”) of all marginal posteriors $f(\bm{x}^{(i)}_{n}|\bm{z}_{1:n})$, $f(b_{\text{bs},n}^{(j)}|\bm{z}_{1:n})$, $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})$, and $p(r_{k,n}|\bm{z}_{1:n})$, for all MT states $\bm{x}^{(i)}_{n}$, BS biases $b_{\text{bs},n}^{(j)}$, and PVAs $k$.

### v2

@@ IV-D Data Association
Data assocation between measurements and PVAs is described in a redundant formulation by the PVA/PA-oriented association vector $\underline{\bm{\mathsfbr{a}}}_{n}^{(i)}\triangleq[\underline{\mathsfbr{a}}_{1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\underline{\mathsfbr{a}}_{J,n}^{(i)}\hskip 2.84526pt\underline{\mathsfbr{a}}_{J+1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\underline{\mathsfbr{a}}_{J+\mathsfbr{K}^{(i)}_{n},n}^{(i)}]^{\mathrm{T}}$ and by the measurement-oriented association vector $\overline{\bm{\mathsfbr{a}}}_{n}^{(i)}\triangleq[\overline{\mathsfbr{a}}_{1,n}^{(i)}\hskip 0.85358pt\cdots\hskip 0.85358pt\overline{\mathsfbr{a}}_{{\mathsfbr{M}}_{n}^{(i)},n}^{(i)}]^{\mathrm{T}}$. See [25, 23, 6, 17] for details.
@@ V-1 Joint Posterior PDF and Factor Graph
Using Bayes’ rule and independence assumptions related to the state-transition PDFs, the prior PDFs, and the likelihood model (for details please see [23, 6, 19]), and for fixed (observed) measurements $\bm{z}_{1:n}$ (the numbers of measurements ${M}^{(i)}_{n}$ are fixed and not random anymore) the joint posterior PDF of $\bm{\mathsfbr{y}}_{0:n}$, $\bm{\mathsfbr{x}}_{0:n}$, $\underline{\bm{\mathsfbr{a}}}_{1:n}$, and $\overline{\bm{\mathsfbr{a}}}_{1:n}$, conditioned on $\bm{z}_{1:n}$ is given by (16), where the individual LHF-related factors are respectively $\underline{q}_{\text{BS}}\big(\bm{x}_{n}^{(i)},{b}^{(j,i)}_{\text{bs},n},\underline{a}^{(i)}_{j,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$, $\underline{q}\big(\underline{\bm{y}}^{(i)}_{k,n},\bm{x}_{n}^{(i)},{b}^{(\underline{j}(k),i)}_{\text{bs},n}\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt,\underline{a}^{(i)}_{k+J,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$ and $\overline{q}\big(\overline{\bm{y}}^{(i)}_{m,n},\bm{x}^{(i)}_{n},\bm{b}^{(i)}_{\text{bs},n}$, $\overline{j}^{(i)}_{m,n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$ that will be discussed next.
@@
$\underline{q}_{\text{BS}}\big(\bm{x}_{n}^{(i)},{b}^{(j,i)}_{\text{bs},n},\underline{a}^{(i)}_{j,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$
$\hskip 28.45274pt\triangleq\begin{cases}\displaystyle\hskip 0.85358pt\frac{p_{\mathrm{d}}f\big({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\big|\hskip 0.85358pt\bm{x}^{(i)}_{n},{b}^{(j,i)}_{\text{bs},n}\big)}{\mu_{\mathrm{fp}}\hskip 0.85358ptf_{\mathrm{fp}}\big({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\big)}\hskip 0.85358pt,&\!\!\hskip-0.85358pt\underline{a}^{(i)}_{j,n}\!=\hskip-0.85358ptm\\[9.95845pt] 1\!-\hskip-0.85358ptp_{\mathrm{d}}\hskip 0.85358pt,&\!\!\hskip-0.85358pt\underline{a}^{(i)}_{j,n}\!=\hskip-0.85358pt0\end{cases}$ (17)
@@
$\underline{q}\big(\underline{\bm{p}}^{(i)}_{k,\text{va}},\underline{r}^{(i)}_{k,n}=1,\bm{x}_{n}^{(i)},{b}^{(\underline{j}(k),i)}_{\text{bs},n},\underline{a}^{(i)}_{k+J,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$
$\triangleq\begin{cases}\displaystyle\hskip 0.85358pt\frac{p_{\mathrm{d}}f\big({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\big|\hskip 0.85358pt\bm{x}^{(i)}_{n},\underline{\bm{p}}^{(i)}_{k,\text{va}}\hskip-0.85358pt,{b}^{(\underline{j}(k),i)}_{\text{bs},n}\big)}{\mu_{\mathrm{fp}}\hskip 0.85358ptf_{\mathrm{fp}}\big({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\big)}\hskip 0.85358pt,&\!\!\hskip-0.85358pt\underline{a}^{(i)}_{k+J,n}\!=\hskip-0.85358ptm\\[9.95845pt] 1\!-\hskip-0.85358ptp_{\mathrm{d}}\hskip 0.85358pt,&\!\!\hskip-0.85358pt\underline{a}^{(i)}_{k+J,n}\!=\hskip-0.85358pt0\end{cases}$ (18)
and $\underline{q}\big(\underline{\bm{x}}^{(i)}_{k,n},\underline{r}^{(i)}_{k,n}=0,\bm{x}_{n}^{(i)},{b}^{(\underline{j}(k),i)}_{\text{bs},n},\underline{a}^{(i)}_{k+J,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)\hskip-0.85358pt\triangleq\hskip-0.85358pt{\delta_{\underline{a}^{(i)}_{k,n}}}$.
The pseudo LHF for new PVA $m$ and MT $i$ is given by
$\overline{q}\big(\overline{\bm{y}}^{(i)}_{m,n},\bm{x}_{n}^{(i)},\bm{b}^{i}_{\text{bs},n},\overline{j}^{(i)}_{m,n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$
$\hskip 14.22636pt\triangleq\begin{cases}0\hskip 0.85358pt,&\hskip-2.84526ptj\!\neq\hskip-0.85358pt\overline{j}^{(i)}_{m,n}\\[2.84526pt] \overline{q}\big(\overline{\bm{y}}^{(i)}_{m,n},\bm{x}_{n}^{(i)},{b}^{(j,i)}_{\text{bs},n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)\hskip 0.85358pt,&\hskip-2.84526ptj\!=\hskip-0.85358pt\overline{j}^{(i)}_{m,n}\end{cases}$ (19)
@@
$\hskip-2.84526pt\overline{q}\big(\overline{\bm{p}}^{(j,i)}_{\text{va},m},\overline{r}^{(j,i)}_{m,n}=1,\bm{x}_{n}^{(i)},{b}^{(j,i)}_{\text{bs},n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)$
$\hskip-2.84526pt\triangleq\begin{cases}0\hskip 0.85358pt,&\hskip-2.84526pt\overline{a}^{(i)}_{m,n}\hskip-0.85358pt\hskip-0.85358pt=\hskip-0.85358ptk\\[2.84526pt] \frac{\mu_{\mathrm{n}}f_{\mathrm{n}}(\overline{\bm{p}}^{(j,i)}_{\text{va},m}\hskip 0.85358pt|\hskip 0.85358pt\bm{x}_{n}^{(i)})f({\bm{z}_{\text{bs}}}^{(i)}_{m,n}\hskip 0.85358pt|\bm{x}_{n}^{(i)},\overline{\bm{p}}^{(j,i)}_{\text{va},m},{b}^{(j,i)}_{\text{bs},n})}{J\,\mu_{\mathrm{fp}}f_{\mathrm{fp}}({\bm{z}_{\text{bs}}}^{(i)}_{m,n})}\hskip 0.85358pt,&\hskip-2.84526pt\overline{a}^{(i)}_{m,n}\hskip-0.85358pt=\hskip-0.85358pt0\end{cases}$ (20)
and $\overline{q}\big(\overline{\bm{p}}^{(j,i)}_{\text{va},m},\overline{r}^{(j,i)}_{m,n}=0,\bm{x}_{n}^{(i)},{b}^{(j,i)}_{\text{bs},n},\overline{a}^{(i)}_{m,n};{\bm{z}_{\text{bs}}}^{(i)}_{n}\big)\hskip-0.85358pt\triangleq\hskip-0.85358ptf_{\text{d}}\big(\overline{\bm{p}}^{(j,i)}_{\text{va},m}\big)$, respectively. Finally, the binary indicator functions that check consistency for any pair $(\underline{a}^{(i)}_{k,n},\overline{a}^{(i)}_{m,n})$ of PVA/PA-oriented and measurement-oriented association variable at time $n$,
@@
$\hskip-5.69054pt\triangleq\begin{cases}0,&\underline{a}^{(i)}_{k,n}=m,\overline{a}^{(i)}_{m,n}\neq k\text{ or }\underline{a}^{(i)}_{k,n}\neq m,\overline{a}^{(i)}_{m,n}=k\\[0.28453pt] 1\hskip 0.85358pt,&\text{otherwise}.\end{cases}$ (21)
@@ V-2 Confirmation of PVAs and State Estimation
We estimate the MT states $\bm{\mathsfbr{x}}^{(i)}_{n}$ (including $\mathsfbr{b}_{\text{mt}\,n}^{(i)}$), the BS biases $\mathsfbr{b}_{\text{bs}\,n}^{(j)}$, and the PVA states $\bm{\mathsfbr{p}}_{\text{va},k}$ from all measurements $\bm{z}_{1:n}$ up to time $n$ using the respective minimum mean-square error (MMSE) estimates [26, Ch. 4] $\hat{{\bm{x}}}^{(i)}_{n}=[\hat{\tilde{\bm{x}}}^{(i)}_{n}\hskip 2.84526pt\hat{\bm{v}}^{(i)}_{n}\hskip 2.84526pt\hat{b}_{\text{mt},n}^{(i)}]$ and $\hat{b}_{\text{bs}\,n}^{(j)}$, which are calculated by the expected values of the respective marginal posterior distributions $f(\bm{x}^{(i)}_{n}|\bm{z}_{1:n})$, $f({b}_{\text{bs},n}^{(j)}|\bm{z}_{1:n})$. Detection of the PVAs $k\!\in\!\{1,\dots,K_{n}\}$ relies on the posterior existence probabilities $p(r_{k,n}\!=\!1|\bm{z}_{1:n})=\int f(\bm{p}_{\text{va},k},r_{k,n}\!=\!1|\bm{z}_{1:n})\,\mathrm{d}\bm{p}_{\text{va},k}$ and the conditional posteriors $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})=f(\bm{p}_{\text{va},k},r_{k,n}\!=\!1|\bm{z}_{1:n})/p(r_{k,n}\!=\!1|\bm{z}_{1:n})$. A PVA is confirmed if $p(r_{k,n}\!=\!1|\bm{z}_{1:n})>p_{\mathrm{cf}}$ [27].
Since introducing new PVAs causes $K_{n}$ to grow indefinitely, a suboptimal pruning step is applied. A PVAs is pruned if $p(r_{k,n}\!=\!1|\bm{z}_{1:n})<p_{\text{pr}}$, except for k=1k\!=\!1 of each BS, which is always retained. Note that before pruining $\underline{j}(k)$ needs to be saved, since (9) becomes invalid. For existing PVAs, we determine the MMSE estimate $\hat{\bm{p}}_{\text{va},k}$ by calculating the expected value of $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})$. Since direct marginalization of (16) is intractable, sequential message passing via the SPA on a factor graph [28] provides efficient approximations (“beliefs”) of all marginal posteriors $f(\bm{x}^{(i)}_{n}|\bm{z}_{1:n})$, $f(b_{\text{bs},n}^{(j)}|\bm{z}_{1:n})$, $f(\bm{p}_{\text{va},k}|r_{k,n}\!=\!1,\bm{z}_{1:n})$, and $p(r_{k,n}|\bm{z}_{1:n})$, for all MT states $\bm{x}^{(i)}_{n}$, BS biases $b_{\text{bs},n}^{(j)}$, and PVAs $k$.